40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Sun Oct 18 19:55:45 2009, rana, Configuration, Electronics, IP POS is back: ND filter gone, new resistors in IMG_0068.JPGIMG_0072.JPG
    Reply  Sun Oct 18 22:06:15 2009, rana, Configuration, Electronics, IP POS is back: ND filter gone, new resistors in Untitled.png
Message ID: 2110     Entry time: Sun Oct 18 19:55:45 2009     Reply to this: 2112
Author: rana 
Type: Configuration 
Category: Electronics 
Subject: IP POS is back: ND filter gone, new resistors in 

Its back in and re-centered. Our next move on IPPOS should be to replace its steering mirror with something bigger and more rigid.

Electronics changes:

20K -> 3.65 K  (R6, R20, R42, R31) (unused)

20K -> 3.65 K  (R7, R21, R32, R43, R11, R24, R35, R46)

If you look in the schematic (D990272), you see that its an AD797 transimpedance stage with a couple of LT1125 stages set to give some switchable gain. It looks like some of these

switches are on and some are not, but I am not sure where it would be controlled from. I've attached a snapshot of one quadrant of the schematic below.

The schematic shows the switches in the so-called 'normally closed' configuration. This is what the switches do with zero volts applied to the control inputs. As the schematic also shows,

just disconnecting the 'switch' inputs cause the switch's control inputs to go high (normally open configuration, i.e. pins 2-3 connected, pin4 open). For the record, the default positions of the IPPOS switches are:

switch1   high

switch2   low

switch3   low

switch4   high

 


** EDIT (Nov 2, 2009): I forgot to attach the before and after images; here they are:

IMG_0068.JPGIMG_0072.JPG

ELOG V3.1.3-